OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value if x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (anonymous):

make a guess, i bet you are right

OpenStudy (anonymous):

@pooja195 @acxbox22

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(3) = 3-4 ?

OpenStudy (anonymous):

@pooja195

OpenStudy (ribhu):

guess its the value of the function at x = 3. which is value of x-4 @ x=3

OpenStudy (anonymous):

-1 @ribhu

OpenStudy (ribhu):

value of f(3) is asked nowhere in the question.

OpenStudy (ribhu):

@hello_hey @pooja195 this is it

OpenStudy (anonymous):

? @ribhu

pooja195 (pooja195):

what is it?

OpenStudy (ribhu):

i explained it no.. 1st option the value of x-4 @ x =3

OpenStudy (ribhu):

@pooja195 i think you must have got it. @hello_hey

OpenStudy (anonymous):

a? @ribhu you are confusing me i need a better explanation

pooja195 (pooja195):

im confused

OpenStudy (anonymous):

same ^^^

OpenStudy (ribhu):

why see lets say we have a function y = f(x)

OpenStudy (anonymous):

@pooja195 may you just help me...

OpenStudy (ribhu):

what is y?? y is the value of f(x) i.e. the function at x.

pooja195 (pooja195):

x=3

OpenStudy (ribhu):

@pooja195 @hello_hey jus be a bit patient and pay attention.

OpenStudy (lizz123):

@satellite73

OpenStudy (ribhu):

@you better tell me which part you didn't understand?? @hello_hey

pooja195 (pooja195):

@hello_hey please stop they are trying to help you

OpenStudy (ribhu):

@pooja195 am i confusing? weren't you getting my explaination?

pooja195 (pooja195):

@ribhu it was a bit unclear :)

OpenStudy (acxbox22):

guys it is a simple explanation plug 3 in for x and solve this gives us no other answer but the first

OpenStudy (ribhu):

question clearly says function is x-4 and there was nowhere mentioned to calculate the value of that function but we have to tell rather what does f(3) represents.

OpenStudy (anonymous):

ok thank you @acxbox22

OpenStudy (acxbox22):

since you arent setting (x-4) equal to 3 you are simply plugging in 3 for x so (3-4) when x is 3

OpenStudy (ribhu):

@acxbox22 there is no need to solve. where it is written? rather tell what is f(3) representing?

OpenStudy (anonymous):

ok, so the answer is a

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

correct :)

OpenStudy (anonymous):

thank you

OpenStudy (ribhu):

so @acxbox22 you didn't explain clearly to @hello_hey

OpenStudy (anonymous):

yes he did @ribhu

OpenStudy (ribhu):

no he told u to calculate value but its not like that.

OpenStudy (anonymous):

well the answer is a right?

OpenStudy (ribhu):

read the answer carefully and explain me its meaning.

OpenStudy (anonymous):

? @ribhu

OpenStudy (acxbox22):

@ribhu you can argue all you want the point is the user understands the explanation and why the answer is what it is after i show them the answer itself using examples

OpenStudy (ribhu):

see it says f(3) represents the value of x-4 (which is the function f(x)) at x = 3. thats what i have been explaining. @hello_hey @acxbox22 @pooja195